A 25.00 mL sample of 0.280 M LiOH is titrated with 0.750 M HI at 25 °C.

initial ph = 13.45

calculate the ph after addition of 5.00 ml titrant is added.
Answer #1

LiOH ------------------------------------------- HI

MB = 0.280M -----------------------------MA = 0.750M

VB   = 25ml -------------------------------VA = 5ml

      M   = MBVB- MAVA/VA + VB            [ MBVB >MAVA]

             = 0.28*25-0.75*5/(5 + 25)

           = 3.25/30

            = 0.1084M

[OH^-] = M   = 0.1084M

POH   = -log[OH^-]
          = -log0.1084

           = 0.965

PH   = 14-POH

        = 14-0.965

          = 13.035 >>>>answer